# Seeing through Satellite Images at Street Views

### v1

@@ 4.5.1 Illumination Modeling
Figure 8: Comparison of image synthesis quality: The second image demonstrates the result from a model trained without GAN loss. Although this model achieves higher image metric scores such as PSNR and SSIM, the synthesized output is visibly poorer in quality. In contrast, the full model, when tested under various illumination inputs, including “Real Illumination”, generates results that, despite achieving different image metric scores, exhibit negligible differences in both image quality and semantic information. This indicates the robustness and consistency of the full model in maintaining visual and semantic fidelity across varying illumination conditions. The first image serves as the ground truth reference.
@@ 4.5.2 Loss Functions
Role of $\mathcal{L}_{\text{opa}}$. The non-sky opacity loss function $\mathcal{L}_{\text{opa}}$ assists in delineating the learning of the 3D ground scene and the 2D sky region within the density field. By employing $\mathcal{L}_{\text{opa}}$, the model is enabled to approximate non-sky areas with opacity values near 1, while sky regions tend towards an opacity of 0. Omitting $\mathcal{L}_{\text{opa}}$ leads to significant generation errors: the 2D sky generator erroneously renders ground features, the satellite and street-view depths rendered are inaccurate compared to the full model, the road region remains static in the generated video, and there is misalignment between the sky and the upper half of the image, as illustrated in the second column of Fig. 6. Without this loss function, the model struggles to differentiate the generation roles of the triplane and sky generators under only image-level supervision, resulting in incorrect region generation.
Role of $\mathcal{L}_{\text{sky}}$. The sky loss $\mathcal{L}_{\text{sky}}$ is crucial for 2D sky image generation. The incorporation of $\mathcal{L}_{\text{sky}}$ leads to notable improvements in realism metrics such as FID and KID. Furthermore, in its absence, the generated sky region in street-view images exhibits more artifacts, as shown in the fourth column of Fig. 6.
@@
Role of illu. + $\mathcal{L}_{\text{opa}}$ + $\mathcal{L}_{\text{sky}}$. In Sat2Density++, we employ illumination modeling, which includes illumination input, the illumination-adaptive NeRF, and the illumination-adaptive 2D sky generator module, to mitigate the effects of illumination on 3D representation learning. Additionally, we utilize $\mathcal{L}_{\text{opa}}$ and $\mathcal{L}_{\text{sky}}$ to segregate the optimization process of 3D scenes from 2D sky regions during training. When these components are removed, the model can be perceived as a vanilla image-conditioned 3D-aware generation model. As shown in Tab. III, performance scores significantly decrease in comparison to the full model, and the generated street view videos are inconsistent and of poor quality, as demonstrated on our project page. This underscores that these training strategies and module designs, specifically tailored to address the SatStreet-view synthesis task’s challenges, are effective in introducing an image-conditioned 3D-aware generation model.
@@
Role of satellite view loss. We apply satellite view reconstruction and GAN losses to align the generated 3D appearance with the input satellite image $I_{sat}$. These losses ensure that the images rendered from the learned radiance field at the satellite viewpoint closely match $I_{sat}$, thereby guaranteeing that the output street-view video remains faithful to the satellite image. As illustrated in Tab. III, omitting the satellite view reconstruction loss results in slightly improved FID and DINO scores. However, without photometric supervision from the input satellite perspective, the learned 3D representation is prone to deviating from the information provided by the satellite view. For instance, Fig. 7 demonstrates that without these losses, the generated zebra crossings exhibit a different style compared to those in the input satellite image. This indicates that incorporating the satellite view losses significantly enhances the faithfulness of the generated ground images and videos to the input satellite image.
@@ 4.6 Applications
To demonstrate the practical utility and robustness of our proposed model, we explore its performance on several downstream applications. These applications serve to validate the effectiveness of our architectural design choices and highlight the potential for real-world impact in various domains.
@@
Figure 11: Selected frames from synthesized video in Seattle City. All the trajectories were filmed sequentially from the red to the blue point. The full videos can be seen on the project page.
@@ 5 Discussion & Limitations
Data challenge. Translating satellite imagery to street views within our conditional 3D generation framework necessitates well-aligned imagery, similar to datasets such as CelebAMask-HQ, AFHQ-cat, and Shapenet-car [49, 50, 51]. However, remote sensing images are not always captured from a nadir perspective, and precise pose calibration of street-view images remains challenging. Moreover, the altitude at which ground images are captured can vary significantly, sometimes from elevated positions like overpasses, which adds to the complexity of positioning within both world and camera coordinate systems—a task crucial for accurately defining our generative space.
Limitations in model design. Our model employs sky region histograms for illumination modeling. While this approach has shown promise, illumination is just one of the many factors that characterize a scene. Incorporating additional scene descriptors such as lighting direction, cast shadows, and albedo may significantly improve synthesis results.
Furthermore, we approximate the sky region as the only area in street-view images that extends beyond the satellite scene. Although this approach has proven effective, it lacks precision: areas extending beyond the satellite coverage might also include distant mountains, buildings, and other elements. Accurately distinguishing and modeling these regions is a direction worthy of future exploration, as it would enhance the accuracy of the learned 3D representations.
@@ 6 Conclusion
In this paper, we studied the task of SatStreet-view synthesis and presented Sat2Density++, a novel approach to synthesize street-view panoramas from a given satellite image as input. Our Sat2Density++ is built upon a deep understanding of SatStreet-view synthesis, devising an effective learning solution by modeling sky regions and street-view illuminations. It approaches the challenging two-view neural field learning problem with a feedforward 3D-aware generation framework for photorealistic street-view panorama synthesis. Extensive experiments on the commonly used suburban scene dataset and the recently proposed urban scene dataset VIGOR demonstrated the effectiveness of our proposed Sat2Density++. To the best of our knowledge, our model is the first that can synthesize multi-view consistent street-view videos from input satellite images, trained without relying on 3D annotations. We hope our proposed method will facilitate 3D representation learning and video generation in outdoor scenes.

### v2

@@ 4.6.1 Sky Illumination Modeling
Figure 9: Comparison of image synthesis quality. The second image, trained without GAN loss, attains higher metric scores but yields visibly poorer results. In contrast, the full model shows negligible differences in visual and semantic quality under varying illumination inputs, demonstrating robustness and consistency. The first image is the ground truth reference.
4.6.2 Illumination Adaptive Tri-plane NeRF
We experimented by replacing the illumination-adaptive tri-plane decoder with a vanilla tri-plane decoder. In this case, the sky illumination input, initially intended to affect both sky region and ground appearances under varying daylight, was limited to adjusting only the sky. Using a vanilla tri-plane decoder resulted in a noticeable decline in image quality compared to the illumination-adaptive decoder, as shown in Tab. III. This occurs because NeRF’s volumetric rendering technique, which generates pixel values based on light-scene interactions, relies on the consistent appearance of multi-view data [37]. Inconsistent lighting prevents parameter optimization.
4.6.3 Dual Branch Design
When the sky branch is removed, the tri-plane NeRF must generate both ground and sky regions simultaneously, presenting two major challenges. First, the tri-plane NeRF can only express a limited spatial range [2], meaning the rendered sky region is given depth and becomes part of the NeRF-generated scene. Consequently, any movement of the camera causes changes in the sky region within the generated panoramic image, which is incorrect because the sky should remain consistent regardless of camera movement, as shown in Fig. 10. Second, it is physically incorrect to represent both the infinitely distant sky and the finite ground regions within a limited 3D space. This increases the model’s learning burden and degrades learning quality, significantly reducing the visual realism of the generated images and videos, as quantitatively demonstrated in Tab. III.
TABLE V: Ablation results for input satellite-view reconstruction on VIGOR-OOD test set.
[…]
w/o sat. view loss 27.5 0.21
Full model 30.0 0.56
Figure 10: Results without the sky branch: includes predicted satellite-view depth and street-view image with depth.
Figure 11: Visualization of results without the sky-region reconstruction loss. We presented three sets of images; in each set, the left image is the generated sky image, and the right is the final generated street-view image.
@@ 4.6.4 Loss Functions
Role of $\mathcal{L}_{\text{opa}}$. The non-sky opacity loss function $\mathcal{L}_{\text{opa}}$ assists in delineating the learning of the 3D ground scene and the 2D sky region within the density field. By employing $\mathcal{L}_{\text{opa}}$, the model is enabled to approximate non-sky areas with opacity values near 1, while sky regions tend towards an opacity of 0. Omitting $\mathcal{L}_{\text{opa}}$ leads to significant generation errors: the 2D sky generator erroneously renders ground features, the satellite and street-view depths rendered are inaccurate compared to the full model, the road region remains static in the generated video, and there is misalignment between the sky and the upper half of the image, as illustrated in the second column of Fig. 7. Without this loss function, the model struggles to differentiate the generation roles of the tri-plane and sky generators under only image-level supervision, resulting in incorrect region generation.
Role of $\mathcal{L}_{\text{sky}}$. The sky loss $\mathcal{L}_{\text{sky}}$ is crucial for 2D sky image generation. The incorporation of $\mathcal{L}_{\text{sky}}$ leads to notable improvements in realism metrics such as FID and KID. As shown in the Fig. 11 and fourth column of Fig. 7, without the proper guidance of sky loss, ground content that should be rendered by the tri-plane NeRF may incorrectly appear in the sky images, such as trees and buildings. This causes the model to fail in accurately distinguishing the generated content between the two branches, leading to unrealistic outcomes.
@@
Role of illu. + $\mathcal{L}_{\text{opa}}$ + $\mathcal{L}_{\text{sky}}$. In Sat2Density++, we employ illumination modeling, which includes illumination input, the illumination-adaptive NeRF, and the illumination-adaptive 2D sky generator module, to mitigate the effects of illumination on 3D representation learning. Additionally, we utilize $\mathcal{L}_{\text{opa}}$ and $\mathcal{L}_{\text{sky}}$ to segregate the optimization process of 3D scenes from 2D sky regions during training. When these components are removed, the model can be perceived as a vanilla image-conditioned 3D-aware generation model. As shown in Tab. III, performance scores significantly decrease in comparison to the full model, and the generated street-view videos are inconsistent and of poor quality, as demonstrated on our project page. This underscores that these training strategies and module designs, specifically tailored to address the SatStreet-view synthesis task’s challenges, are effective in introducing an image-conditioned 3D-aware generation model.
@@
Role of satellite-view loss. We apply satellite-view reconstruction and GAN losses to align the generated 3D appearance with the input satellite image $I_{sat}$. These losses ensure that the images rendered from the learned radiance field at the satellite viewpoint closely match $I_{sat}$, thereby guaranteeing that the output street-view video remains faithful to the satellite image. As illustrated in Tab. III, omitting the satellite-view reconstruction loss results in slightly improved FID and DINO scores. However, without photometric supervision from the input satellite perspective, the learned 3D representation is prone to deviating from the information provided by the satellite view. Quantitatively, Tab. V shows that removing the satellite-view loss leads to a notable drop in satellite image reconstruction quality on the VIGOR-OOD test set (PSNR decreases from 30.0 to 27.5, and SSIM from 0.56 to 0.21). In Fig. 8, the predicted satellite images show deterioration in the absence of satellite supervision, lacking consistency with the input satellite imagery, such as in crosswalks, trees, and building appearances, and the generated zebra crossings in the predicted street-view images exhibit a different style compared to those in the input satellite image. This indicates that incorporating the satellite-view losses significantly enhances the faithfulness of the generated ground images and videos to the input satellite image.
@@ 4.7 Applications
To demonstrate the practical utility and robustness of our proposed model, we explore downstream applications. These applications serve to validate the effectiveness of our design and highlight the potential for real-world impact.
@@
Figure 14: Selected frames from the synthesized video in Seattle City. All the trajectories were filmed sequentially from the red to the blue point. The full videos can be seen on the project page.
@@ 5 Discussion & Limitations
Data challenge. Translating satellite imagery to street views within our conditional 3D generation framework necessitates well-aligned imagery, similar to datasets such as CelebAMask-HQ, AFHQ-cat, and Shapenet-car [23, 4, 3]. However, satellite images are not always captured from a nadir perspective, and precise pose calibration of street-view images remains challenging. Moreover, the altitude of ground images can vary considerably, for example, when captured from elevated positions such as overpasses, further complicating their placement in both world and camera coordinate systems, a step that is critical for accurately defining the generative space.
Use sky region color histogram to represent the sky illumination. We acknowledge the potential of enriching illumination inputs with data such as sun position, cloud distribution, or atmospheric radiation to better handle unknown conditions and improve video quality. However, this would deviate from our focus and introduce substantial complexity. Daylight estimation from outdoor imagery is a challenging and ill-posed problem [22, 70, 12, 20], and relying on existing models may add noise while requiring extra effort for our synthesis framework to interpret. Given the demonstrated effectiveness of the simple sky-region histogram, we defer incorporating richer illumination cues to future work.
3D space for generation. We approximate the sky region as the only area in street-view images that extends beyond the satellite scene. Although this approach has proven effective, it lacks precision: areas extending beyond the satellite coverage might also include distant mountains, buildings, and other elements. Accurately distinguishing and modeling these regions is a direction worthy of future exploration, as it would enhance the accuracy of the learned 3D representations.
@@ 6 Conclusion
In this paper, we studied the task of SatStreet-view synthesis and presented Sat2Density++, a novel approach to synthesize street-view panorama video from a given satellite image as input. Our Sat2Density++ is built upon a deep understanding of SatStreet-view synthesis, devising an effective learning solution by modeling sky regions and street-view illuminations. It approaches the challenging two-view neural field learning problem with a feedforward 3D-aware generation framework for photorealistic street-view panorama synthesis. Extensive experiments on the commonly used suburban scene dataset and the recently proposed urban scene dataset VIGOR demonstrated the effectiveness of our proposed Sat2Density++. To the best of our knowledge, our model is the first that can synthesize multi-view consistent street-view videos from input satellite images, trained without relying on 3D annotations. We hope our proposed method will facilitate 3D representation learning and video generation in outdoor scenes.